how to calculate ratio surface atoms to volume for 1 meter gold sphere
Answer #1

1 meter gold sphere means that the diameter of the sphere is 1 meter.

Diameter = 1 m

radius = 0.5 m

surface area of sphere = 4\pir2

volume of sphere = 4/3 \pir3

ratio of surface area to volume =

\frac{4\pi r^{2}}{\frac{4}{3}\pi r^{3}} = \frac{3}{r}

On solving, we get 3/r.

r = 0.5 m

So, ratio of surface area to volume = 3/0.5 = 6 m-1